OpenStudy (anonymous):

What is the exact diameter of a circular rug with a circumference of seventy-two feet?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\pi D = 72\) solve \(D\)

ganeshie8 (ganeshie8):

since they want exact D, simply divide pi both sides

ganeshie8 (ganeshie8):

\(\pi D = 72\) divide pi both sides \( \large D = \frac{72}{\pi} \)

OpenStudy (anonymous):

22.9

ganeshie8 (ganeshie8):

dont approximate like that

ganeshie8 (ganeshie8):

they want exact answer, that means they want \(\large \frac{72}{\pi}\)

ganeshie8 (ganeshie8):

DONOT APPRXIMATE when they ask for EXACT ANSWER remember this for rest for ur course time

OpenStudy (anonymous):

ok so 72/ π

ganeshie8 (ganeshie8):

yup !